# High-low analysis and small cap decoupling over non-Archimedean local fields

### v1

@@ 1 Introduction
For $p,q\geq 1$ and $R,\beta$ as above, we will write $D_{p,q}^{\mathbb{K}}(R;\beta)$ for the infimal constant such that, for any Schwartz-Bruhat function111i.e. a finite linear combination of indicators of metric balls. $f:\mathbb{K}^{2}\to\mathbb{C}$ with Fourier support contained in $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$, we have
$\|f\|_{L^{p}(\mathbb{K}^{2})}^{p}\leq D_{p,q}^{\mathbb{K}}(R;\beta)\left(\sum_{\gamma\in\Gamma_{\beta}(R^{-1})}\|f_{\gamma}\|_{L^{p}(\mathbb{K}^{2})}^{q}\right)^{p/q}.$
@@
Let $p,q\geq 1$ satisfy $\frac{3}{p}+\frac{1}{q}\leq 1$, $R\geq\mathbf{p}^{32}$, and $\beta\in[\frac{1}{2},1]$. Then the small cap decoupling constant satisfies
$D_{p,q}^{\mathbb{K}}(R;\beta)\leq\frac{10^{6}}{(\log\mathbf{p})^{16}}\mathbf{p}^{12}(\log R)^{16+6\beta^{-1}}\Big{(}R^{\beta(p-\frac{p}{q}-1)-1}+R^{p\beta(\frac{1}{2}-\frac{1}{q})}\Big{)}.$ (1.1)
Here $\mathbf{p}=\mathbf{p}_{\mathbb{K}}$ is defined as:
$\mathbf{p}=\begin{cases}p&\mathbb{K}\text{ extends }\mathbb{Q}_{p}\text{ or }\mathbb{F}_{p}(\!(t)\!),\,p>2,\\ 2^{d}&\mathbb{K}\text{ extends }\mathbb{Q}_{2},\,[\mathbb{K}:\mathbb{Q}_{2}]=d.\end{cases}$
The study of non-Archimedean decouplings was initiated in [6], where a bilinear variant of an inequality of the form $D_{p,2}^{\mathbb{Q}_{q}}(R;\frac{1}{2})\lesssim_{q,\varepsilon}(\log R)^{2p+\varepsilon}$ was established for $q>2$ to achieve good discrete restriction estimates for the parabola. It was continued in [15], where that author observed many of the features of non-Archimedean analysis that made it particularly appropriate for the setting of decoupling. In [13], the current author and Lin generalized the decoupling theorem for the moment curve to the $q$-adic setting for each $q>n$ (with $n$ the ambient dimension); the result there may be extended to any local field of characteristic either 00 or greater than $n$.
@@
Over the real numbers, small cap decouplings were introduced in [3], and the estimate $D_{p,p}^{\mathbb{R}}(R;\beta)\lesssim_{\varepsilon}R^{\beta(\frac{p}{2}-1)+\varepsilon}$ was proved (with the reasonable interpretation of $D_{p,q}^{\mathbb{R}}$). The work of [5] established the real version of the estimate we will show, which is a superlevel set estimate implying sharp bounds on the $D_{p,q}^{-}(R;\beta)$ in the regime $\frac{1}{q}+\frac{3}{p}\leq 1$. Substantial work has also generalized the idea of small cap decouplings to other manifolds, such as that contained in [7] and [8].
@@
Let $f:\mathbb{K}^{2}\to\mathbb{C}$ be Schwartz-Bruhat with Fourier support in $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$. Then, for any $\alpha>0$,
$\alpha^{4}\mu\big{(}\{x:|f(x)|>\alpha\}\big{)}\leq 13\mathbf{p}^{12}(\log R)^{10}\sum_{\begin{subarray}{c}s\in\mathbf{p}^{\mathbb{Z}}\\ R^{-1/2}\leq s\leq 1\end{subarray}}\sum_{\begin{subarray}{c}\tau\\ \mathrm{diam}(\tau)=s\end{subarray}}\sum_{U\in\mathcal{G}_{\tau}}\mu(U)\left(\fint_{U}\sum_{\theta\subseteq\tau}|f_{\theta}|^{2}\right)^{2}.$
Here we use the following notation: each $U\in\mathcal{G}_{\tau}$ is a rectangle of dimensions $R\times sR$, with long edge in the direction of the normal vector to $\mathbb{P}^{1}$ at the center of $\tau$, centered at 00; the set $\mathcal{G}_{\tau}$ is the subset of the standard tiling of $\mathbb{K}^{2}$ by such rectangles for which the following holds:
@@
Here #$\#\tau$ denotes the number of $\tau$ of a particular length for which $f_{\tau}\not\equiv 0$. We also write $\theta$ for a generic cap on $\mathbb{P}^{1}$ of dimensions $R^{-1/2}\times R^{-1}$.
@@
We will restrict our attention to cases (b), (c), i.e. the non-Archimedean local fields of characteristic 00, resp. $p>2$. The reader will generally benefit from imagining the cases $\mathbb{Q}_{p},\mathbb{F}_{p}(\!(t)\!)$ with $p>2$.
We also mention the utility in tracking dependence on $\mathbf{p}$ in the above theorems. Over the reals, Fourier-analytic methods for counting solutions to Diophantine equations frequently entail a subpolynomial loss in the diameter of the variable set. For instance, the Bourgain-Demeter-Guth resolution [2] of the main conjecture of Vinogradov’s mean value theorem shows the particular result
@@
[…]
@@
Here $\delta_{p}(\mathcal{A})=\min\{|n-m|_{p}:n\neq m\in\mathcal{A}\}$, where $|\cdot|_{p}$ is the usual $p$-adic norm. In particular, if $p$ does not divide any of the differences $n-m$ (say, if $p>\max(|n|:n\in\mathcal{A})$), then the subpolynomial factor trivializes. However, there are corresponding losses in the choice of prime. Thus, judicious tracking of the dependence on $p$, together with number-theoretic considerations, may reduce the dependence on features of $\mathcal{A}$ other than its cardinality. See [18] for another approach to the same problem.
@@
Lastly, we discuss the special role of the prime $p=2$ in the above; particularly, why we have excluded the fields $\mathbb{F}_{2^{d}}(\!(t)\!)$, and why $\mathbf{p}$ is much larger for extensions of $\mathbb{Q}_{2}$. In the case of characteristic $2$, say $\mathbb{K}=\mathbb{F}_{2^{d}}(\!(t)\!)$, by the Frobenius identity $(x+y)^{2}=x^{2}+y^{2}$ we see that $\mathbb{P}^{1}$ is contained in a linear subspace of $\mathbb{K}^{2}$, when the latter is regarded a vector space over $\mathbb{F}_{p^{d}}$. In particular, linear equations among the first powers of frequency variables imply corresponding equations among the second powers. A straightforward computation (working first over even integer exponents, then interpolating and comparing with the trivial Cauchy-Schwarz bounds) supplies the identity $D_{p,2}^{\mathbb{F}_{2^{d}}(\!(t)\!)}(R;\frac{1}{2})=R^{\frac{1}{2}(\frac{p}{2}-1)}$, for any $R\in 2^{d\mathbb{N}}$ and $p\geq 2$. One may compare with the proof of the local bilinear restriction estimate Theorem (2.4) below to see the effect of the size of $2$ in $\mathbb{K}$, when $2\neq 0$ is small.
@@ 1.1 Brief overview of method
Here the sets $\delta\theta^{(k)}$ are understood as follows: if $\theta$ is the cap about 00 for simplicity, that is, $\theta=B(0,R^{-1/2})\times B(0,R^{-1})$, such that in particular $\theta=\delta\theta$ (in our non-Archimedean setting), then for $0\leq k\leq N$ we write $\delta\theta^{(k)}$ for
@@
(writing of course $\mathcal{P}_{A}$ for the Fourier projection onto a set $A$), valid whenever $k>0$ and the $\tau$ have diameter $d(\tau)\geq R^{\frac{k}{2N}-1}$; and the pointwise estimate
@@
valid whenever $k>0$ and the $\tau$ have diameter $d(\tau)\geq R^{-\frac{|k|}{N}}$. (1.3) and (1.4) are known as the high and low lemmas, respectively; see Lemmas (2.2) and (2.1) below. By repeatedly applying estimates of the form (1.3) and (1.4), together with the usual $L^{4}$ Córdoba-Fefferman square function estimate, we find that for each $k\geq\ell$ we have that

### v2

@@ 1 Introduction
For $p,q\geq 1$ and $R,\beta$ as above, we will write $D_{p,q}^{\mathbb{K}}(R;\beta)$ for the infimal constant such that, for any Schwartz–Bruhat function111i.e. a finite linear combination of indicators of metric balls. $f:\mathbb{K}^{2}\to\mathbb{C}$ with Fourier support contained in $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$, we have
$\|f\|_{L^{r}(\mathbb{K}^{2})}^{r}\leq D_{r,q}^{\mathbb{K}}(R;\beta)\left(\sum_{\gamma\in\Gamma_{\beta}(R^{-1})}\|f_{\gamma}\|_{L^{r}(\mathbb{K}^{2})}^{q}\right)^{r/q}.$
@@
Let $r,q\geq 1$ satisfy $\frac{3}{r}+\frac{1}{q}\leq 1$, $R\geq\mathbf{p}^{32}$, and $\beta\in[\frac{1}{2},1]$. Then the small cap decoupling constant satisfies
$D_{r,q}^{\mathbb{K}}(R;\beta)\leq 2\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{11}}(\log R)^{17+6\beta^{-1}}\Big(R^{\beta(r-\frac{r}{q}-1)-1}+R^{\beta(\frac{r}{2}-\frac{r}{q})}\Big).$ (1.1)
Here $\mathbf{q}=p^{f}=\#\mathbf{k}$ is the cardinality of the residue field of $\mathbb{K}$, and $\mathbf{p}=\mathbf{p}_{\mathbb{K}}$ is defined as:
$\mathbf{p}=\begin{cases}\mathbf{q}&\mathbb{K}\text{ extends }\mathbb{Q}_{p}\text{ or }\mathbb{F}_{p}(\!(t)\!),\,p>2\,,\\ 2^{d}&\mathbb{K}\text{ extends }\mathbb{Q}_{2},\,[\mathbb{K}:\mathbb{Q}_{2}]=d.\end{cases}$ (1.2)
Note in particular that, when $\mathrm{char}(\mathbf{k})\neq 2$, the loss in the size $\mathbf{q}=\mathbf{p}$ of the residue field is reduced to $\frac{\mathbf{p}^{7}}{(\log\mathbf{p})^{11}}$. The actual inequality we obtain in the proof of Theorem (1.1) in Section (3) is slightly more refined; here we have stated a simplified version.
The study of non-Archimedean decouplings was initiated in [6], where a bilinear variant of an inequality of the form $D_{r,2}^{\mathbb{Q}_{q}}(R;\frac{1}{2})\lesssim_{r,\varepsilon}(\log R)^{2r+\varepsilon}$ was established for $q>2$ to achieve good discrete restriction estimates for the parabola. It was continued in [16], where Li observed many of the features of non-Archimedean analysis that made it particularly appropriate for the setting of decoupling. In [14], the current author and Lin generalized the decoupling theorem for the moment curve to the $q$-adic setting for each $q>n$ (with $n$ the ambient dimension); the result there may be extended to any local field of characteristic either 0 or greater than $n$.
@@
Over the real numbers, small cap decouplings were introduced in [3], and the estimate $D_{r,r}^{\mathbb{R}}(R;\beta)\lesssim_{\varepsilon}R^{\beta(\frac{r}{2}-1)+\varepsilon}$ was proved (with the natural interpretation of $D_{r,q}^{\mathbb{R}}$). The work of [5] established the real version of the estimate we will show, which is a superlevel set estimate implying sharp bounds on the $D_{r,q}^{\square}(R;\beta)$ in the regime $\frac{1}{q}+\frac{3}{r}\leq 1$. Substantial work has also generalized the idea of small cap decouplings to other manifolds, such as that contained in [7] and [8].
@@
Let $f:\mathbb{K}^{2}\to\mathbb{C}$ be Schwartz–Bruhat with Fourier support in $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$. Then, for any $\alpha>0$,
$\alpha^{4}\mu\big(\{x:|f(x)|>\alpha\}\big)\leq\frac{1}{50}\mathbf{q}^{-3}\mathbf{p}^{10}(\log_{\mathbf{p}}R)^{10}\sum_{\begin{subarray}{c}s\in\mathbf{p}^{\mathbb{Z}}\\ R^{-1/2}\leq s\leq 1\end{subarray}}\sum_{\begin{subarray}{c}\tau\\ \mathrm{diam}(\tau)=s\end{subarray}}\sum_{U\in\mathcal{G}_{\tau}}\mu(U)\left(\fint_{U}\sum_{\theta\subseteq\tau}|f_{\theta}|^{2}\right)^{2}.$
Here we use the following notation: each $U\in\mathcal{G}_{\tau}$ is a rectangle of dimensions $R\times sR$, with long edge in the direction of the normal vector to $\mathbb{P}^{1}$ at the center of $\tau$, centered at 0; the set $\mathcal{G}_{\tau}$ is the subset of the standard tiling of $\mathbb{K}^{2}$ by such rectangles for which the following holds:
@@
Here $\tau$ denotes the number of $\tau$ of a particular length for which $f_{\tau}\not\equiv 0$. We also write $\theta$ for a generic cap on $\mathbb{P}^{1}$ of dimensions $R^{-1/2}\times R^{-1}$. The choice of $\mathcal{G}_{\tau}$ in subsection (2.2) is more technical, but (by the “pruning inequalities”) will include strictly fewer envelopes $U$ than in the statement of Theorem (1.2), hence is sufficient.
@@
We will restrict our attention to cases (b), (c), i.e. the non-Archimedean local fields of characteristic 0, resp. $p>2$. The reader will generally benefit from imagining the cases $\mathbb{Q}_{p},\mathbb{F}_{p}(\!(t)\!)$ with $p>2$.
We also mention the utility in tracking dependence on $\mathbf{p}$ in the above theorems. Over the reals, Fourier-analytic methods for counting solutions to Diophantine equations frequently entail a subpolynomial loss in the diameter of the variable set. For instance, the Bourgain–Demeter–Guth resolution [2] of the main conjecture of Vinogradov’s mean value theorem shows the particular result
@@
whenever $\mathcal{A}\subseteq\mathbb{Z}$ has $\mathcal{A}=A$. Here we use the usual notation of
@@
Here $\delta_{p}(\mathcal{A})=\min\{|n-m|_{p}:n\neq m\in\mathcal{A}\}$, where $|\cdot|_{p}$ is the usual $p$-adic norm. In particular, if $p$ does not divide any of the differences nmn-m (say, if $p>\max(|n|:n\in\mathcal{A})$), then the subpolynomial factor trivializes. However, there are corresponding losses in the choice of prime. Thus, judicious tracking of the dependence on $p$, together with number-theoretic considerations, may reduce the dependence on features of $\mathcal{A}$ other than its cardinality. See [19] for another approach to the same problem.
@@
Lastly, we discuss the special role of the prime $p=2$ in the above; particularly, why we have excluded the fields $\mathbb{F}_{2^{d}}(\!(t)\!)$, and why $\mathbf{p}$ is much larger for extensions of $\mathbb{Q}_{2}$. In the case of characteristic $2$, say $\mathbb{K}=\mathbb{F}_{2^{d}}(\!(t)\!)$, by the Frobenius identity $(x+y)^{2}=x^{2}+y^{2}$ we see that $\mathbb{P}^{1}$ is contained in a linear subspace of $\mathbb{K}^{2}$, when the latter is regarded a vector space over $\mathbb{F}_{p^{d}}$. In particular, linear equations among the first powers of frequency variables imply corresponding equations among the second powers. A straightforward computation (working first over even integer exponents, then interpolating and comparing with the trivial Cauchy–Schwarz bounds) supplies the identity $D_{p,2}^{\mathbb{F}_{2^{d}}(\!(t)\!)}(R;\frac{1}{2})=R^{\frac{1}{2}(\frac{p}{2}-1)}$, for any $R\in 2^{d\mathbb{N}}$ and $p\geq 2$. One may compare with the proof of the local bilinear restriction estimate Theorem (2.4) below to see the effect of the size of $2$ in $\mathbb{K}$, when $2\neq 0$ is small.
1.1 Applications
To motivate the study of non-Archimedean decoupling theorems, we present some applications of Theorem (1.1). We emphasize that all of the integral estimates obtained in this section are real, i.e. integrals of functions over subsets of $\mathbb{R}^{n}$. In the terminal section (4), we will supply the needed transference procedure.
Our first application is a variant of the well-known fact that small cap decoupling theorems imply “short mean-value estimates,” i.e. mean value estimates along certain subsets of the fundamental domain. We first demonstrated this consequence in [13]. For comparison, we first state the usual result.
Theorem 1.3 (c.f. Corollary 1 of [5]).
Let $N\in\mathbb{N}$ and $\sigma\in[0,1].$ Let $a_{1},\ldots,a_{N}\in\mathbb{C}$ be arbitrary. Let $1\leq r,q<\infty$ satisfy $\frac{3}{r}+\frac{1}{q}\leq 1$. Then we have
$\fint_{[0,1]\times[0,N^{\sigma-1}]}\Bigg|\sum_{n=1}^{N}a_{n}e\Big(nx+n^{2}t\Big)\Bigg|^{r}dxdt\lesssim_{\varepsilon}N^{\varepsilon}\big(N^{(r-\frac{r}{q}-1)-1-\sigma}+N^{\frac{r}{2}-\frac{r}{q}}\big)\Bigg(\sum_{n=1}^{N}|a_{n}|^{q}\Bigg)^{r/q}.$ (1.4)
Our variant replaces the “short” domain $0,1]\times[0,N^{\sigma-1}]$ with a “sparse” domain.
Definition 1.4.
Let $\sigma,\gamma\in[0,1]$ be two parameters. Let $N\in\mathbb{N}$ be arbitrary. The $(\sigma,\gamma)$-sparse domain at scale $N$ is the set
$A(N,\sigma;\gamma):=\bigcup_{j=0}^{N^{(1+\sigma)(1-\gamma)}-1}[0,1]\times[jN^{-\sigma(1-\gamma)},jN^{-\sigma(1-\gamma)}+N^{\gamma(1+\sigma)-2}].$ (1.5)
Remark 1.5.
If $\gamma=1$, the union includes only one term which is simply $0,1]\times[0,N^{\sigma-1}]$.
Alternately, if $\gamma=0$, then $A(N,\sigma;1)$ is essentially the vertical $N^{-1}$-neighborhood of the maximal progression of horizontal lines $0,1]\times\{0\}$, $0,1]\times\{N^{-\sigma}\},[0,1]\times\{2N^{-\sigma}\},\ldots$, in the range $0\leq t<1$.
As $\gamma$ ranges from 0 to $1$, the number of thickened lines decreases and the thickness increases.
Remark 1.6.
It always holds that $\mathcal{L}^{2}\big(A(N,\sigma;\gamma)\big)=N^{\sigma-1}$ (as long as $N^{(1+\sigma)(1-\gamma)}$ is an integer). Thus, the $A(N,\sigma;\gamma)$ may be viewed as “cousins” of the short domain $0,1]\times[0,N^{\sigma-1}]$.
With this definition, we may state our first application.
Theorem 1.7.
Let $N\in\mathbb{N}$ and $\sigma,\gamma\in[0,1]$. Suppose that $N,N^{(1+\sigma)(1-\gamma)}$ are both powers of a common odd prime $p$. Let $a_{1},\ldots,a_{N}\in\mathbb{C}$ be arbitrary. Let $1\leq r,q<\infty$ satisfy $\frac{3}{r}+\frac{1}{q}\leq 1$. Then we have
$\fint_{A(N,\sigma;\gamma)}\Bigg|\sum_{n=1}^{N}a_{n}e\Big(nx+n^{2}t\Big)\Bigg|^{r}dxdt\lesssim(\log N)^{O(1)}\big(N^{(r-\frac{r}{q}-1)-1-\sigma}+N^{\frac{r}{2}-\frac{r}{q}}\big)\Bigg(\sum_{n=1}^{N}|a_{n}|^{q}\Bigg)^{r/q}.$ (1.6)
If $\gamma=0$, the $(\log N)^{O(1)}$ term may be refined to $(\log N)^{23+6\sigma}$.
Thus, the usual “short” (or “small ball”) mean value estimates (1.4) fit into a family of identical mean value estimates (1.6).
Remark 1.8.
By a variant of the tensor product argument that will be used to combine real and $p$-adic decoupling for the proof of Theorem (1.7), it will be possible to somewhat relax the condition that $N,N^{(1+\sigma)(1-\gamma)}$ are powers of a common odd prime. Indeed, by taking tensors over distinct $\mathbb{Q}_{p}$ with $p$ ranging over a small number of odd primes, we can allow $N,N^{(1+\sigma)(1-\gamma)}$ to have prime factorization composed of a small number of small odd primes. We do not pursue this here.
The proof of Theorem (1.7) involves three steps. First, we prove a variant of Theorem (1.3) working purely over the $p$-adics for an arbitrary choice of $p$. Second, we demonstrate a “transference procedure,” carrying integrals of certain $p$-adic exponential sums to integrals of real exponential sums over a special domain; at this point, we have handled the $\gamma=0$ case of Theorem (1.7). Lastly, to deal with $0<\gamma<1$, we regard the desired exponential sum as a sum of tensor products of real and $p$-adic exponential sums, and apply the general fact that decoupling theorems tensorize (see e.g. Lemma 5.5 of [14]) to estimate the exponential sum along sets resembling a product of $p$-adic intervals with real intervals.
A related estimate, Prop. (3.1), which is obtained on the way to proving Theorem (1.1), implies the following superlevel set estimate for exponential sums. This is the “non-Archimedean” variant of Theorem 4 from [5].
Theorem 1.9.
Fix $\sigma\in[0,1]$, $\varepsilon>0$, and $p\equiv 3$ (mod 4)($\mathrm${mod\ 4}). Let $N\in\mathbb{N}$ be such that $N,N^{\sigma}$ are both powers of $p$ and such that $p\leq D_{\varepsilon}N^{\varepsilon/20}$. Let $S\subseteq\{1,\ldots,N\}$. Let $\{a_{j}\}_{j\in S}$ be a family of complex numbers with $|a_{j}|\leq 1$.
Define the distribution functions:
$\lambda_{k}(S)=\max_{0\leq a<p^{k}}\#\big\{j\in S:j\equiv a\quad(\mathrm{mod}\,\,p^{k})\big\}.$
Let $f$ be the exponential sum
$f(x,t)=\sum_{j\in S}a_{j}e\Big(\frac{j}{N}x+\frac{j^{2}}{N^{2}}t\Big).$
Let $1\leq\alpha\leq N$ and $U_{\alpha}=\{(x,t)\in[0,1]^{2}:|f(x,t)|\geq\alpha\}$. Then we have the superlevel set estimate
$\frac{\mathcal{L}^{2}\big(U_{\alpha}\cap A(N,\sigma;0)\big)}{\mathcal{L}^{2}\big(A(N,\sigma;0)\big)}\lesssim_{\varepsilon}p^{7}\alpha^{-4}D_{\varepsilon}N^{\varepsilon}\max_{\frac{1+\sigma}{2}\log_{p}N\leq k\leq\log_{p}N}\big(\lambda_{k}(S)\cdot\lambda_{(1+\sigma)\log_{p}N-k}(S)\big)\sum_{j\in S}|a_{j}|^{4}.$
Our third application involves a great many possible examples; we select the simplest one, which we believe adequately illustrates the general theme. In contrast to the two results above, this involves quoting the decoupling theorem over a field extension of some $\mathbb{Q}_{p}$.
For $\sigma\in[0,1]$ and $N\in\mathbb{N}$, we write
$A^{(2)}(N,\sigma)=\bigcup_{j,k=0}^{N^{(1+\sigma)}-1}[0,1]^{2}\times[jN^{-\sigma-1},jN^{-\sigma-1}+N^{-2}]\times[kN^{-\sigma-1},kN^{-\sigma-1}+N^{-2}].$
Theorem 1.10.
Let $N\in\mathbb{N}$ and $\sigma\in[0,1]$. Let $\{a_{nm}\}_{1\leq n,m\leq N}$ be arbitrary complex numbers. Let $1\leq r,q<\infty$ satisfy $\frac{3}{r}+\frac{1}{q}\leq 1$. Then we have
$\begin{split}\fint_{A^{(2)}(N,\sigma)}\Bigg|\sum_{n,m=1}^{N}a_{nm}&e\Big(nx-my+z(n^{2}-m^{2})-2wnm\Big)\Bigg|^{r}dxdydzdw\\ &\lesssim(\log N)^{23+6\sigma}\big(N^{(r-\frac{r}{q}-1)-1-\sigma}+N^{\frac{r}{2}-\frac{r}{q}}\big)\left(\sum_{n,m=1}^{N}|a_{nm}|^{q}\right)^{r/q}.\end{split}$ (1.7)
Remark 1.11.
The phase function may be substituted with many equivalents, e.g.
$nx+my+z(n^{2}-m^{2})+2wmn.$
Theorem (1.10) follows from essentially the same analysis as was used in Theorem (1.7), with the $\mathbb{Q}_{p}$ in the latter substituted for $\mathbb{K}=\mathbb{Q}_{p}\big[\sqrt{-1}\big]$ subject to $p\equiv 3\,\,(\mathrm{mod}\,4)$.
The proofs of Theorems (1.7), (1.9), and (1.10) will be postponed to the end in Section (4).
@@ 1.2 Brief overview of method
Here the sets $\delta\theta^{(k)}$ are understood as follows: if $\theta$ is the cap about 0 for simplicity, that is, $\theta=B(0,R^{-1/2})\times B(0,R^{-1})$, such that in particular $\theta=\delta\theta$ (in our non-Archimedean setting), then for $0\leq k\leq N$ we write $\delta\theta^{(k)}$ for
@@
(writing $\mathcal{P}_{A}$ for the Fourier projection onto a set $A$), valid whenever $k>0$ and the $\tau$ have diameter $d(\tau)\geq R^{\frac{k}{2N}-1}$; and the pointwise estimate
@@
valid whenever $k>0$ and the $\tau$ have diameter $d(\tau)\geq R^{-\frac{|k|}{N}}$. (1.8) and (1.9) are known as the high and low lemmas, respectively; see Lemmas (2.2) and (2.1) below. By repeatedly applying estimates of the form (1.8) and (1.9), together with the usual $L^{4}$ Córdoba–Fefferman square function estimate, we find that for each $k\geq\ell$ we have that